OpenStudy (anonymous):

Help! Medals awarded! For the sequence an=4n+5 Find the sum of the first 9 terms

OpenStudy (anonymous):

i need some serious help from anyone willing to try

OpenStudy (ikram002p):

ok its easy locate n from 1,...,9 now u need to find a1,.....,a9 i'm gonna solve a1 & a2 and u try wid the rest of them a1=4(1)+5=4+5=1 a2=4(2)+5=8+5=13 got it or still confused ??

OpenStudy (magbak):

@werty00

OpenStudy (anonymous):

I'm still kinda confused..

OpenStudy (ikram002p):

sry i thought u wanna the first 9 turms not the sum , ok to find the sum \[\sum_{1}^{9}4n+5\] =\[ 4 \sum_{1}^{9}n + (5\times (9-1))\] do u know this formula ??

OpenStudy (anonymous):

432

OpenStudy (anonymous):

the formula is SUM = n/2 * (2a1 + (n-1) * d) = 9/2 * (2*1 + (9-1) * 12)

OpenStudy (ikram002p):

i think the answer 184 at the end